OpenStudy (anonymous):

Evaluate 23-7 divided by (1/3+1/4)

OpenStudy (anonymous):

the question is not clear. Use the equation toolbox and write down your question. You can type in the equation use the Equation button just underneath this area.

OpenStudy (anonymous):

\[23-7\div(1/3+1/4)\]

OpenStudy (anonymous):

thanks for telling me about that...I didn't know how to input my equation.

OpenStudy (anonymous):

np. so, first evaluate 1/3 +1/4

OpenStudy (anonymous):

i find the LCM of both of the them right? If so it's12

OpenStudy (anonymous):

right, so 1/3 = 4/12 and 1/4 = 3/12

OpenStudy (anonymous):

that's what I got. yay! okay now we add them it becomes (7/12)

OpenStudy (anonymous):

okay great! so what is 7 / 7 / 12?

OpenStudy (anonymous):

dividing right 7/7 = 1 and 7/12 = 1.71

OpenStudy (anonymous):

huh? no. it is \[7 \div (7/12)\]

OpenStudy (anonymous):

you have to be patient with me. I haven't done math in 8 years. :(

OpenStudy (anonymous):

\[7\times12/7\] 7 and 7 cancel out. so the answer is 12.

OpenStudy (anonymous):

23-12 = 11 is your final answer.

OpenStudy (anonymous):

oh...I see what I did wrong. duh!

OpenStudy (anonymous):

heh. \[a \div b = a \times 1/b\]

OpenStudy (anonymous):

So if I was going to do this problem..... (0.95-0.68)x6. I would do whats inside the parenthesis first and complete it before I move on to multiply. That rule goes for all with parenthesis......

OpenStudy (anonymous):

yes. you should always follow order of operations http://en.wikipedia.org/wiki/Order_of_operations#The_standard_order_of_operations

OpenStudy (anonymous):

PEMDAS Please, Excuse, My, Dear, Aunt, Sally~~~

OpenStudy (anonymous):

P - Parenthesis, E= exponents, M = Multiplication, D= division, A = addition, S = subtraction ^_^ and + always start computing from left to right

OpenStudy (anonymous):

also, BODMAS.

OpenStudy (anonymous):

it's weird though, but I don't have an aunt called sally o_o lol, but anyhow, hope this sets things right ^_^

OpenStudy (anonymous):

LOL! what does that stand for?

OpenStudy (anonymous):

yeah. I am still learning how this all works again. We are working on equations and decimals, and word problems. I am having such a hard time with this stuff.

OpenStudy (anonymous):

nah, you'll get through with it in no time ^_^

OpenStudy (anonymous):

B Brackets first O Orders (ie Powers and Square Roots, etc.) DM Division and Multiplication (left-to-right) AS Addition and Subtraction (left-to-right)

OpenStudy (anonymous):

lol, my dear aunt sally statement is better :D

OpenStudy (anonymous):

I always used aunt sally

OpenStudy (anonymous):

oh well, I was taught bodmas. its all the same.

OpenStudy (anonymous):

likewise LOL, I know dhat ^_^

OpenStudy (anonymous):

sigh, I miss school lol

OpenStudy (anonymous):

lol, yeah. I would be so smart at it now!

OpenStudy (anonymous):

yep! >_< and I'd get 100s and As all over the place

OpenStudy (anonymous):

true. :)

OpenStudy (anonymous):

^_^